A 1200 kg car is coasting on a horizontal road with a speed of 18 m/s . After passing over an unpaved, sandy stretch 35.0 m long, car's speed has decreased to 12 m/s .

Part A Was the net work done on the car positive, negative, or zero?

Part B Find the magnitude of the average net force on the car in the sandy section of the road. Express your answer to two significant figures and include the appropriate units.
Answer #1

as, according to work energy principal the net work done in changing a motion of particle is equal to the change in its kinetic energy.

so,

W = KEf - KEi

where W is the work done, KEf and KEi are the final and initial kinetic energy of the car.

W = (1/2)×m×(vf^2 - vi^2)

where m is the mass of the car, vf and vi are the final and initial velocity of the car.

W = (1/2)(1200 kg)[(12 m/s)^2 - (18 m/s)^2] = -10800 J

(A) now, as we can see the net work done on the car is negative.

(B) F(avg) = W/d = (-10800 J)/(35 m) = -3085.71 N